# ActAlign: Zero-Shot Fine-Grained Video Classification via Language-Guided Sequence Alignment

### v1

@@ Abstract
We address the task of zero-shot fine-grained video classification, where no video examples or temporal annotations are available for unseen action classes. While contrastive vision–language models such as SigLIP demonstrate strong open-set recognition via mean-pooled image–text similarity, they fail to capture the temporal structure critical for distinguishing fine-grained activities. We introduce ActAlign, a zero-shot framework that formulates video classification as sequence alignment. For each class, a large language model generates an ordered sub-action sequence, which is aligned with video frames using Dynamic Time Warping (DTW) in a shared embedding space. Without any video–text supervision or fine-tuning, ActAlign achieves 30.5% accuracy on the extremely challenging ActionAtlas benchmark, where human accuracy is only 61.6%. ActAlign outperforms billion-parameter video–language models while using approximately 8$\times$ less parameters. These results demonstrate that structured language priors, combined with classical alignment techniques, offer a scalable and general approach to unlocking the open-set recognition potential of vision–language models for fine-grained video understanding.
Codehttps://github.com/aghdamamir/ActAlign
@@ 1 Introduction
Understanding fine-grained human activities in video—such as distinguishing a hook shot from a layup in basketball, or recognizing tactical formations in football—requires parsing subtle, temporally ordered visual cues across frames. These actions unfold in structured sequences of sub-events and are often nearly indistinguishable from one another in appearance. In contrast to general activities like swimming, which can often be inferred from a single frame showing a person in water, fine-grained recognition demands attention to temporally extended object interactions, spatial relations, and high-level intent. As such, models must not only understand what is present in a video but also when and how key sub-actions occur. This requires accurately aligning the temporal progression of sub-actions with each fine-grained activity to ensure a correct prediction, as shown in Figure 1.
At the same time, contrastive vision–language models such as CLIP (Radford et al. 2021) and SigLIP (Zhai et al. 2023) have demonstrated impressive open-set fine-grained recognition in static image domains by training on massive image–text pairs. These contrastive models enable zero-shot classification using natural language prompts and have been widely adopted for downstream recognition tasks. However, extending these capabilities to video understanding introduces new challenges and requires temporal modeling. Existing methods that adapt CLIP-style models to video recognition either average frame-level features (Rasheed et al. 2023; Zohra et al. 2025)—ignoring temporal structure—or fine-tune on target datasets (Wang, Xing, and Liu 2022; Ni et al. 2022; Kim et al. 2024; Wang et al. 2024a), sacrificing generalization and open-set recognition. In both cases, the fine-grained temporal semantics of actions are lost or diluted.
Recent video–language architectures and instruction-tuned LLM-based systems such as Video-LLaMA (Zhang, Li, and Bing 2023), VideoChat (KunChang Li 2023) mPLUG-Owl (Ye et al. 2023), Qwen2-VL (Wang et al. 2024b), and DeepSeek-JanusPro (Chen et al. 2025) enable open-ended, dialog-style video understanding through heavy instruction tuning, but they are not tailored for fine-grained video recognition.
Meanwhile, textual grounding (image–text alignment) remains a central challenge in interactive video-language models, especially for open-set and fine-grained video recognition. Dynamic Time Warping (DTW) (Vintsyuk 1968), a classical algorithm for aligning temporally mismatched sequences, has seen renewed interest through differentiable variants (Chang et al. 2019; Dogan et al. 2018) designed for supervised image–text temporal alignment. Yet, these methods rely on annotated transcripts or example support videos, making them impractical for zero-shot recognition. Likewise, approaches using part-level or attribute-level supervision (Wu et al. 2023; Zhu et al. 2024) offer fine-grained cues but lack the ability to model the temporal structure between language-defined actions and visual content.
@@
Figure 1: ActAlign significantly improves zero-shot fine-grained action recognition by modeling them as structured language sequences. By aligning sub-action descriptions with video frames (green vs. red paths), we achieve more accurate predictions without requiring any video-text training data.
Figure 2: Our pipeline uses an LLM to generate structured sub-action sequences for each fine-grained candidate class in ActionAtlas (Salehi et al. 2024), forming SubActionAtlas. This structured representation enables alignment with video content for zero-shot recognition tasks.
In this work, we introduce ActAlign, a novel framework that brings the open-set generalization power of image–text models to fine-grained video recognition through language-guided subaction alignment in a truly zero-shot setting. Rather than tuning a model for a specific domain or collapsing the video into a static representation, ActAlign operates in a training-free setting: for each unseen action class, we use a large language model (LLM) to generate a structured sequence of temporal sub-actions that semantically define the class. (see Figure 2. Then, using the pretrained SigLIP model (Zhai et al. 2023) to extract frame-wise visual and subaction features, we align the video sequence with the LLM-generated subaction script via Dynamic Time Warping (DTW).(see Figure 1) This allows us to compute a soft alignment score between different action classes that respects both content and temporal ordering, enabling fine-grained classification in a truly zero-shot manner.
To rigorously evaluate our approach, we use ActionAtlas (Salehi et al. 2024), the most challenging fine-grained video recognition benchmark to our knowledge, with a human accuracy ceiling of 61.64%. It comprises sports footage paired with extremely fine-grained candidate tactics that demand close attention to subtle visual cues for correct classification. We construct SubActionAtlas, by decomposing each candidate action into a sequence of sub-actions via LLM prompting (Figure 2), providing structured templates for alignment.
@@
  • We introduce a novel framework for fine-grained video recognition that models each action as a general, structured temporal sequence of sub-actions derived solely from action names—without access to videos or transcripts.
  • We propose ActAlign, a novel zero-shot framework that applies the open-set generalization strength of contrastive image–text models to the challenging task of fine-grained and open-set video classification, without requiring any video–text supervision.
  • We show that ActAlign consistently outperforms prior zero-shot and CLIP-based baselines, and even exceeds billion-parameter video–language models on challenging fine-grained video classification benchmarks.
@@ 2 Related Work
2.1 Contrastive Image–Language Models
@@
Contrastive vision–language models such as CLIP (Radford et al. 2021), SigLip (Zhai et al. 2023), and ALIGN (Jia et al. 2021) learn joint image–text embeddings from large-scale image–caption pairs, enabling strong open-set recognition without task-specific training. These models serve as pretrained backbones for downstream tasks, including visual question answering and reasoning (Li, Baldridge, and Hoi 2021; Li et al. 2022; Tsimpoukelli et al. 2021), image captioning and generation (Mokady, Hertz, and Bermano 2021; Wang et al. 2021b), and few-shot and zero-shot classification (Zhou et al. 2021; Khattak et al. 2025). Notably, SigLip (Zhai et al. 2023) recently enhanced CLIP’s architecture and leveraged a larger training corpus to enhance in image–text matching (Zhai et al. 2023). However, such models lack any temporal structure for processing video inputs. This frame-level abstraction limits their open-set classification ability for video inputs.
2.2 Video–Language Models
Video-Language Modeling
Initial efforts to extend vision–language models to video focused on pairing visual frames with corresponding narrations or transcripts. MIL-NCE (Miech et al. 2020) introduced multi-instance contrastive learning to align videos and narrations in uncurated instructional datasets. Later approaches leveraged transformer architectures to model temporal sequences. VideoBERT (Sun et al. 2019) and ActBERT (Zhu, Xu, and Yang 2020) applied masked language modeling over sequences of video frames and transcripts, To improve computational efficiency, ClipBERT (Lei et al. 2021) proposed sparse frame sampling for end-to-end video–text alignment.
Interactive Video Language Models
Recent work integrates large language models (LLMs) with visual encoders to support open-ended video understanding tasks such as captioning and dialogue. Systems like Video-LLaMA (Zhang, Li, and Bing 2023), mPLUG-Owl (Ye et al. 2023), and VideoChat (KunChang Li 2023) combine pretrained visual backbones with chat-centric LLMs to generate detailed responses and spatiotemporal reasoning in conversational settings. For example, VideoChat enables multi-turn dialogue grounded in video content, while Video-LLaMA augments an LLM with temporal and audio query transformers. These systems excel at descriptive and interactive tasks but require extensive instruction tuning (Zhang, Li, and Bing 2023; KunChang Li 2023) and are not optimized for fine-grained video recognition.
CLIP for Video Classification
Pretrained CLIP-style Image-Language models exhibit strong open-set image recognition capability, motivating adaptations to utilize their power for video recognition tasks. These adaptations include temporal modules and special prompting strategies. ActionCLIP (Wang, Xing, and Liu 2022) introduced a “pre-train, prompt, and fine-tune” strategy, augmenting CLIP with handcrafted label prompts and training on action datasets. X-CLIP (Ni et al. 2022) extended CLIP with temporal attention layers to process frame sequences. Other adaptations inject motion-awareness via learnable visual prompts or adapters (Ju et al. 2022; Lin et al. 2022). While these approaches show strong performance on closed-set benchmarks like Kinetics, their effectiveness relies on dataset-specific fine-tuning. As shown in ViFi-CLIP (Rasheed et al. 2023), such specialization often harms zero-shot generalization by overfitting to seen classes.
2.3 Video-Text Sequence Alignment
Classic Alignment Methods
Aligning sequences of video frames to textual descriptions has long been a challenge, especially under weak supervision where frame-level labels are unavailable. Classic approaches often rely on Dynamic Time Warping (DTW) or similar sequence alignment algorithms. For example, Bojanowski et al. (Bojanowski et al. 2015) proposed aligning videos to ordered scripts by enforcing the temporal order of events, without precise timestamps. Other works adopt the Connectionist Temporal Classification (CTC) loss—originally developed for speech recognition—for action segmentation in video. Richard et al. (Richard, Kuehne, and Gall 2018) and Huang et al. (Huang et al. 2016) used CTC and Viterbi decoding to align video frames with a sequence of action labels, learning to segment actions without frame-level supervision. However, these approaches typically assume access to ground-truth transcripts (i.e., the precise ordered list of sub-actions) for every training video.
Dynamic Time Warping (DTW) for Video Classification
Beyond supervision regimes, alignment algorithms themselves have evolved to improve flexibility and learning capacity. Differentiable variants of DTW, such as D3TW (Chang et al. 2019), introduced continuous relaxations that enable end-to-end gradient-based optimization under ordering constraints. NeuMATCH (Dogan et al. 2018) proposed a neural alignment model using recurrent moving windows to align long video sequences to textual inputs. More recently, graph-based models such as the Dynamic Graph Warping Transformer (Wang et al. 2021a) have incorporated structured reasoning and global constraints to improve alignment accuracy. OTAM (Cao et al. 2020) utilizes DTW for few-shot video classification. These methods still rely on task-specific supervision such as paired video–text exemplars or support videos from the target classes.
@@ 2.3 Video-Text Sequence Alignment
Unlike prior work that either discards temporal structure, fine-tunes away generality, or assumes support data for alignment, ActAlign bridges vision and language via sequence-to-sequence matching—paving the way for generalizable, interpretable, and temporally grounded video understanding in zero-shot settings. Our method targets zero-shot video classification of fine-grained actions by leveraging language as a structured latent representation aligned to visual content, without requiring video-text training data or fine-tuning.
@@ 3.2 Our Method
3.3 Preliminary Subaction Generation by LLM
In domains requiring extremely fine-grained distinctions— for example, differentiating between tactical plays in sports footage—the high-level action class name $c_{m}$ alone often lack sufficient discriminatory power. To reveal subtle inter-class differences, it is critical to decompose each action into an ordered sequence of granular, trackable sub-actions. We therefore define a mapping $\mathcal{P}:\mathcal{Y}\to\mathcal{S}$, where $\mathcal{S}$ is the space of all possible sub-action scripts, and for each fine-grained action $c_{m}$, $\mathcal{P}(c_{m})=S_{m}=[s_{m,1},\dots,s_{m,K_{m}}]$ denotes the LLM-generated sequence of sub-actions.
To instantiate $\mathcal{P}$, we employ a pretrained large-language model (GPT-4o) via carefully engineered natural language prompts. Given the set of candidate action class names ${c_{1},\dots,c_{M}}$, our prompt instructs the model to:
@@ 3.3 Preliminary Subaction Generation by LLM
    Decompose it into a varriable-sized sequence of self-sufficient, observable steps (sub-actions) that are semantically coherent and temporally ordered.
@@
    Return the list $[s_{m,1},\dots,s_{m,K_{m}}]$ for each $c_{m}$ in a consistent and structured format.
We then parse the model’s output to form $S_{m}$. By leveraging the LLM’s extensive world knowledge and linguistic priors, this process yields high-quality subaction sequence without any manual annotation or video examples. These scripts serve as the semantic reference signals for subsequent temporal alignment and classification.
@@
We find that terse sub-action descriptions (e.g., “drive forward”) are too ambiguous that can refer to several sports, making DTW alignment ineffective. In contrast, context-rich prompts (e.g., “drive forward to the rim in basketball”) yield semantically grounded sub-actions that align well with video frames and support effective sequence matching.
@@ 3.4 Visual and Semantic Feature Encoding
Let $\phi_{v}:\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{d}$ denote the vision encoder. For video $V_{i}=\{\mathbf{v}i^{t}\}_{t=1}^{T_{i}}$, we compute frame-level embeddings as shown in Eq. 4:
@@
Let $\phi_{t}:\mathcal{T}\to\mathbb{R}^{d}$ be the text encoder. For each class $c_{m}$ and its subaction sequence $S_{m}$, we embed each step as shown in Eq. 5:
@@
is a valid cosine similarity measure between frame $\mathbf{v}_{i}^{t}$ and sub-action $s_{m,k}$ of class $c_{m}$ (see Eq. 6). This cross-modality similarity forms the basis for alignment in the next step.

### v2

@@ Abstract
We address the task of zero-shot video classification for extremely fine-grained actions (e.g., Windmill Dunk in basketball), where no video examples or temporal annotations are available for unseen classes. While image–language models (e.g., CLIP, SigLIP) show strong open-set recognition, they lack temporal modeling needed for video understanding. We propose ActAlign, a truly zero-shot, training-free method that formulates video classification as a sequence alignment problem, preserving the generalization strength of pretrained image–language models. For each class, a large language model (LLM) generates an ordered sequence of sub-actions, which we align with video frames using Dynamic Time Warping (DTW) in a shared embedding space. Without any video–text supervision or fine-tuning, ActAlign achieves 30.5% accuracy on ActionAtlas—the most diverse benchmark of fine-grained actions across multiple sports—where human performance is only 61.6%. ActAlign outperforms billion-parameter video–language models while using $\sim$8$\times$ fewer parameters. Our approach is model-agnostic and domain-general, demonstrating that structured language priors combined with classical alignment methods can unlock the open-set recognition potential of image–language models for fine-grained video understanding.
Codehttps://github.com/amir-aghdam/ActAlign
@@ 1 Introduction
Understanding fine-grained human activities in video—such as distinguishing a hook shot from a layup in basketball, or recognizing tactical formations in football—requires parsing subtle, temporally ordered visual cues across frames. These fine-grained actions unfold in structured sequences of sub-actions and are often nearly indistinguishable from one another in appearance. In contrast to general activities like swimming, which can often be inferred from a single frame showing a person in water, fine-grained recognition demands attention to temporally extended object interactions, spatial relations, and high-level intent. As such, models must not only understand what is present in a video but also when and how key sub-actions occur. This requires accurately aligning the temporal progression of sub-actions with each fine-grained action to ensure a correct prediction, as shown in Figure 1. While, their high-level label is barely recognized by action recognition models (e.g. such as Roundhouse kick in Kickboxing).
@@
Figure 1: ActAlign improves zero-shot fine-grained action recognition by modeling them as structured language sequences. By aligning sub-action descriptions with video frames (green vs. red paths), we achieve more accurate predictions without requiring any video-text training data.
At the same time, contrastive vision–language models such as CLIP (Radford et al. 2021) and SigLIP (Zhai et al. 2023) have demonstrated impressive open-set fine-grained recognition in image domains by training on massive image–text pairs. These contrastive models enable zero-shot classification using natural language prompts and have been widely adopted for downstream recognition tasks. However, extending these capabilities to video understanding introduces new challenges and requires temporal modeling. Existing methods that adapt CLIP-style models to video recognition either average frame-level features (Rasheed et al. 2023; Zohra et al. 2025)—ignoring temporal structure—or fine-tune on target datasets (Wang, Xing, and Liu 2022; Ni et al. 2022; Kim et al. 2024; Wang et al. 2024a), sacrificing generalization and open-set recognition. In both cases, the fine-grained temporal semantics of actions are lost or diluted.
Recent video–language architectures and instruction-tuned LLM-based systems such as Video-LLaMA (Zhang, Li, and Bing 2023), VideoChat (Li et al. 2023) mPLUG-Owl (Ye et al. 2023), Qwen2-VL (Wang et al. 2024b), and DeepSeek-JanusPro (Chen et al. 2025) enable open-ended, dialog-style video understanding through heavy instruction tuning, but they are not tailored for fine-grained video recognition.
Meanwhile, textual grounding (image–text alignment) remains a central challenge in interactive video-language models, especially for open-set and fine-grained video recognition. Dynamic Time Warping (DTW) (Vintsyuk 1968), a classical algorithm for aligning temporally mismatched sequences, has seen renewed interest through differentiable variants (Dogan et al. 2018; Chang et al. 2019) designed for supervised image–text temporal alignment. Yet, these methods rely either on annotated transcripts or example support videos or originally proposed for supervised training, making them impractical for zero-shot recognition. Likewise, approaches using part-level or attribute-level supervision (Wu et al. 2023; Zhu et al. 2024) offer fine-grained cues but lack the ability to model the temporal structure between language-defined actions and visual content.
In this work, we introduce ActAlign, a novel framework that brings the open-set generalization power of image–text models to video classification of extremely fine-grained actions through language-guided sub-action alignment in a truly zero-shot setting. Rather than tuning a model for a specific domain or collapsing the video into a static representation, ActAlign operates in a training-free setting: for each unseen action class, we use a large language model (LLM) to generate a structured sequence of temporal sub-actions that semantically define the class. Then, using the pretrained SigLIP model (Zhai et al. 2023) to extract frame-wise visual and subaction features, we align the frame sequence with the LLM-generated sub-action script via Dynamic Time Warping (DTW) (see Figure 1). This allows us to compute a soft alignment score between different action classes that respects both content and temporal ordering, enabling fine-grained classification in a truly zero-shot manner.
@@
  • We introduce a novel approach for zero-shot fine-grained video recognition that models each action as a general, structured temporal sequence of sub-actions derived solely from action names—without access to videos or transcripts.
  • We propose ActAlign, a domain-general, model-agnostic and zero-shot framework that applies the open-set generalization strength of image–text models to the challenging task of fine-grained and zero-shot video classification by reformulating the task to sequence matching without requiring any video–text supervision.
  • We demonstrate that ActAlign surpasses zero-shot and CLIP-based model baselines, outperforming even billion-parameter video–language models on the most challenging and diverse benchmark to our knowledge, ActionAtlas—where human-level accuracy caps at 61.64%.
@@ 2 Related Work
2.1 Action Recognition
Action recognition has been extensively studied, with early work leveraging hand-crafted features and classic classifiers (Dalal and Triggs 2005; Jain, Jégou, and Bouthemy 2013; Wang and Schmid 2013). The field has since evolved to deep learning approaches (Carreira and Zisserman 2017; Simonyan and Zisserman 2014; Tran et al. 2018; Wang et al. 2016), including CNNs, RNNs, and more recently, Transformer-based models (Arnab et al. 2021; Bertasius, Wang, and Torresani 2021; Kim et al. 2025; Yoshida et al. 2025), which model spatio-temporal dynamics more effectively. Fine-grained action recognition remains challenging due to subtle inter-class variations over time. Recent methods tackle this under supervised (Leong et al. 2022; Sun et al. 2022), semi-supervised (Li, He, and Xu 2022; Dave, Rizve, and Shah 2025), and few-shot (Tang et al. 2023; Hong et al. 2021; Wang et al. 2021a) settings to mitigate annotation costs. Yet, they remain reliant on some level of fine-grained labeling and are constrained to specific domains.
2.2 Vision-Language Models
Image–Language Models
@@ 2.2 Vision-Language Models
Foundational image–language models such as CLIP (Radford et al. 2021), SigLip (Zhai et al. 2023), and ALIGN (Jia et al. 2021) learn joint image–text embeddings from large-scale image–caption pairs. Such modelling enables strong open-set recognition without task-specific supervision. These models are widely adopted as pretrained backbones for downstream tasks, including visual question answering (Li, Baldridge, and Hoi 2021; Tsimpoukelli et al. 2021; Li et al. 2022b; Steiner et al. 2024), image captioning and generation (Mokady, Hertz, and Bermano 2021; Wang et al. 2021b; Mahjourian and Nguyen 2025), and few-/zero-shot classification (Zhou et al. 2021; Khattak et al. 2025). However, they lack temporal modeling, which limits their open-set recognition capability to video inputs.
Video–Language Models
Early efforts in video modeling focused on self-supervised pretraining. Models like ActBERT(Zhu and Yang 2020) and VideoBERT(Sun et al. 2019) applied masked language modeling to video frames. This enabled better transfer to video–language tasks through finetuning on paired video–text data (Miech et al. 2020; Lei et al. 2021; Xu et al. 2021; Feichtenhofer, Fan, and He 2022; Li et al. 2022a). Recent progress in LLM reasoning has driven their integration with visual encoders, enabling open-ended video understanding via user prompts. These models (Ye et al. 2023; Li et al. 2023; Zhang et al. 2024; Wang et al. 2024b), such as Video-LLaMA (Zhang, Li, and Bing 2023), combine pretrained visual backbones with chat-centric LLMs to generate spatio-temporal reasoning in conversational settings. While these models excel at open-ended question answering, they require extensive instruction tuning and are not optimized for fine-grained video recognition.
2.3 Sequence Alignment
Sequence alignment is a long-standing problem in fields such as biology and signal processing. A widely used technique for aligning temporally misaligned signals is Dynamic Time Warping (DTW)(Berndt and Clifford 1994; Keogh and Pazzani 2000; Rath and Manmatha 2003), which computes the optimal warping path between two ordered sequences. Early work extended DTW to video tasks(Huang et al. 2016; Richard, Kuehne, and Gall 2018). For instance, a seminal approach (Bojanowski et al. 2015) proposed aligning videos to ordered scripts by enforcing the temporal order of events. With the rise of deep learning, DTW has been adapted into differentiable forms (Dogan et al. 2018; Chang et al. 2019), enabling greater flexibility and improved learning capacity. A notable example is OTAM (Cao et al. 2020), which leverages DTW for representation learning in few-shot video classification by aligning each video to its support set. However, these methods often assume access to ground-truth transcripts for each video or rely on task-specific supervision, such as paired video–text exemplars.
@@ 2.3 Sequence Alignment
In contrast to prior methods that discard temporal structure, rely on task-specific supervision, or require support data for alignment, ActAlign introduces a language-guided sequence alignment framework for zero-shot fine-grained video classification.
@@ 3.2 Our Method
3.3 Preliminary Sub-action Generation by LLM
In domains requiring fine-grained distinctions—such as differentiating tactical plays in sports—the high-level action class name $c_{m}$ often lacks sufficient discriminatory power. To address this, we define a mapping $\mathcal{P}:\mathcal{Y}\to\mathcal{S}$, where $\mathcal{S}$ denotes the space of ordered sub-action scripts. For each class $c_{m}\in\mathcal{Y}$, we generate a sequence $S_{m}=[s_{m,1},\dots,s_{m,K_{m}}]=\mathcal{P}(c_{m})$, where each $s_{m,k}$ describes a trackable and temporally ordered sub-action.
We instantiate $\mathcal{P}$ using GPT-4o via carefully constructed natural language prompts. Given a set of candidate classes $\{c_{1},\dots,c_{M}\}$, the LLM is instructed to:
@@ 3.3 Preliminary Sub-action Generation by LLM
    Decompose each $c_{m}$ into a variable- or fixed-length sequence of semantically coherent sub-actions.
@@
    Return the list $[s_{m,1},\dots,s_{m,K_{m}}]$ in a consistent, structured format.
By leveraging the LLM’s extensive prior knowledge, we obtain sub-action sequences without any manual labels or video supervision. These sequences serve as semantic reference signals for temporal alignment.
@@
Terse descriptions (e.g., “drive forward”) are often ambiguous and may refer to multiple sports, which hampers DTW alignment due to poor textual grounding. In contrast, context-rich prompts (e.g., “player drives forward to the rim”) produce more discriminative sub-actions that are reliably grounded with visual evidence.
Context Augmentation
We enrich each sub-action by embedding it into a contextualized template: This is a video of doing <action name> in <sport domain> with <subaction>. This substantially improves semantic clarity and further disambiguate sub-actions.
@@ 3.4 Visual and Semantic Feature Encoding
Let $\phi_{v}:\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{d}$ denote the vision encoder. For video $V_{i}=\{\mathbf{v}i^{t}\}_{t=1}^{T_{i}}$, we compute frame-level embeddings:
@@
Let $\phi_{t}:\mathcal{T}\to\mathbb{R}^{d}$ be the text encoder. For each class $c_{m}$ and its subaction sequence $S_{m}$, we embed each step:
@@
is a cosine similarity measure between frame $\mathbf{v}_{i}^{t}$ and sub-action $s_{m,k}$ of class $c_{m}$ (see Eq. 6). This cross-modality similarity forms the basis for alignment in the next step.